# Compound AI Systems Optimization: A Survey of Methods, Challenges, and Future Directions

### v1

@@ 1 Introduction
The community has witnessed a new generation of AI systems centered on large language models (LLMs), incorporating several sophisticated components such as simulators, code interpreters, web search tools, and retrieval-augmented generation (RAG) modules. These systems have shown remarkable capabilities across domains and typically outperform standalone LLMs. For instance, LLMs communicating with symbolic solvers can tackle Olympiad-level math problems Trinh et al. (2024), integrate with search engines and code interpreters to match human programmers’ performance Li et al. (2022); Yang et al. (2024), and when coupled with knowledge graphs, drive biological materials discovery Ghafarollahi and Buehler (2024).
Figure 1: The proposed 2$\times$2 taxonomy spans Structural Flexibility (y-axis) and Learning Signals (x-axis). Representative methods for each quadrant along with their key designs and potential drawbacks.
@@
However, even with mature toolkits that streamline the design process of compound AI systems, such as, LangChain Chase (2022) and LlamaIndex Liu (2022), substantial human intervention remains essential to tailor these systems toward targeted downstream applications Xia et al. (2024); Zhang et al. (2024c), often involving trial-and-error tuning prompt templates and system pipelines based on heuristics. This limitation has motivated recent efforts to develop principled, automated methods for end-to-end AI system optimization. Yet, the operational schemes of these approaches diverge largely depending on whether they allow modifications to system topology and on how they carry learning signals. Moreover, the field still lacks standardized terminology or a cohesive conceptual framework, making some articles difficult for newcomers to navigate Cheng et al. (2024); Khattab et al. (2023). Despite existing surveys Lin et al. (2024); Liu et al. (2025) have provided helpful frameworks, they focus on optimization driven by natural language, overlook critical schemes that permit updates to system topology and do not cover the most recent advances.
To address these gaps, this study introduces four key dimensions for examining existing methods and, drawing on two of these dimensions, constructs a 2$\times$2 taxonomy (Fig. 1) that holistically covers 26 representative works. Setting a reasonable scope for our survey, we excluded related but not directly relevant work. These include prompt optimization techniques for a single LLM Zhou et al. (2022b); Pryzant et al. (2023); Yang et al. (2023); Das et al. (2024); Guo et al. (2025), systems built for task-specific applications without optimization Zhuge et al. (2023); Hong et al. (2023), and papers that do not explicitly frame their problem as system optimization Zhou et al. (2022a); Sordoni et al. (2023). The proposed taxonomy is inherently adaptable to future research, and we hope it will enable developers of compound AI systems to gain a high-level overview before delving into technical details, while allowing researchers proposing novel algorithms to better situate and compare their contributions.
The remainder of this paper is organized as follows: Sec. 2 provides the necessary background and formalization for compound AI systems and their optimization. Sec. 3.1 discusses the four key dimensions, followed by Sec. 3.2, which presents a detailed review of surveyed papers based on the proposed 2$\times$2 taxonomy. Sec. 4 identifies the main challenges facing current methods and their corresponding future directions, and Sec. 5 concludes with a summary of our contributions.
@@ 2.2 Formal Definitions
Specifically, we use $\Phi=(G,\mathcal{F})$ to denote a compound AI system that inputs a user query $q\in\mathcal{Q}$ and outputs a response (answer) $a\in\mathcal{A}$. In particular, $G=(V,E)$ is a directed graph and =$\mathcal{F}=\{f_{i}\}_{i=1}^{|V|}$ is a set of operations (e.g., LLM forward pass, RAG step, external tool calling). Each $f_{i}$ is attached to node $v_{i}$, producing
@@ 3.2 Representative Methods
[…]
@@
Serving as pioneers in the field, these methods demonstrate that learning from text is possible not only for single AI models but also for systems composed of discrete modules. However, their successes are supported mainly by empirical results without theoretical grounding. Moreover, the common reliance on proprietary LLMs to generate NL feedback incurs high API costs.
@@
Rather than relying on textual signals, methods in this category use numerical signals to update node parameters. DSPy Khattab et al. (2023) provides a Python library [^5] http://dspy.ai featuring declarative programming modules for designing and optimizing $\Phi$. As a method that learns directly from raw system metrics, it introduces a suite of rejection-sampling–based routines (Bootstrap-*) for generating high-quality in-context demonstrations informed by corresponding system performance. Users may optionally fine-tune LLM weights ($\theta_{i,N}$) on the collected demonstrations via BootstrapFinetune procedure. MIPRO Opsahl-Ong et al. (2024) advances by jointly optimizing demonstrations and instructions. Specifically, it employs a Bayesian surrogate model to maintain and update posterior distributions over instruction–demonstration configurations, favoring those that yield high performance. BetterTogether Soylu et al. (2024) extends by alternating between LLM prompt and weight optimization, enabling LLMs to iteratively teach themselves and outperform single-strategy baselines.
@@
These methods introduce novel ways to leverage system performance signals, effectively mitigating the imprecision of natural language. However, excessive human-designed rules may limit generalizability. Additionally, fine-tuning each LLM in $\Phi$ incurs substantial GPU resource demands.
@@
Methods discussed previously tune only node parameters within a predefined system topology. To overcome this constraints, methods in this category leverage NL Feedback to jointly optimize both node parameters and overall system structure. Concurrent with TextGrad Yuksekgonul et al. (2025), Agent Symbolic Learning Zhou et al. (2024) designs optimizers with three components: PromptOptimizerToolOptimizer, and PipelineOptimizer, making it goes beyond node tuning because the latter two components support tool creation as well as node addition, deletion, and movement. Through a series of analytical experiments, MASS Zhou et al. (2025) demonstrates that optimizing LLM prompts $\{\theta_{i,T}\}$ can deliver performance gains in compound AI systems more easily than exploring the topology $(V,E)$. Building on this insight, MASS devises a three-stage framework in which prompt optimization precedes topology search.
@@
Methods in this category demonstrate how novel algorithms can leverage the power of proprietary LLMs for direct workflow design. However, effective engagement with these meta LLMs demands high token consumption Hu et al. (2024) or multiple LLM inferences Su et al. (2025). Moreover, evaluations on less powerful open-source models remain scarce.
@@
Completing our review, this passage describes methods that employ Numerical Signals to refine $\Phi$ without posing fixed topology constraints. DyLAN Liu et al. (2024) and GPTSwarm Zhuge et al. (2024) are methods in this category that do not involve model fine-tuning. DyLAN models multiturn debate as a temporal feed-forward network in which agents prompted with distinct roles are unrolled across layers. The system is optimized through pruning unhelpful agents and adaptively connect the surviving agents between consecutive layers. Rule-based algorithms are employed to score each agent’s importance, positioning DyLAN within the category that leverages raw system metrics as learning signals. GPTSwarm models $\Phi$ as a hierarchical framework composed of nodes ($v_{i}\in V$), agents (graphs that link nodes), and swarms (composite graphs that interconnect multiple agents), then employs an optimization procedure to refine the edge connections among agents. A parameterized probabilistic distribution $D_{\theta}$ is introduced to govern the connectivity within the swarm graph, which is then optimized using a gradient-ascent variant of the REINFORCE Williams (1992) algorithm, placing GPTSwarm within the class of RL methods.
Before discussing the remaining methods, we note that they operate in a query-adaptive manner, i.e., instantiating a new $\Phi$ for each $q_{i}$. This operating scheme contrasts with methods discussed so far in which a single task-specific $\Phi$, once optimized, is used for all $q_{i}$. Similar to ADAS Hu et al. (2024) and AFlow Zhang et al. (2024a), these methods leverage a meta LLM to generate optimal system designs. However, rather than relying solely on inference, they fine-tune the meta LLM using constructed training datasets or reward functions to enable effective pipeline generation, given the limited prior knowledge about compound AI system design within LLMs
Starting from the SFT category, MAS-GPT Ye et al. (2025) first constructs a query pool from open-source queries across various domains, and a MAS pool by manually implementing over 40 common system designs in Python code. Through evaluation, selection, and refinement of query–MAS pairs, MAS-GPT curates a training dataset that ensures consistency, minimizing ambiguity when fine-tuning the meta LLM.
@@
Recent trends in this category frame compound system optimization as a meta LLM fine-tuning problem, thereby sidestepping the non-differentiability of modules within $\Phi$. However, substantial effort is still required to curate high-quality training data, and the lack of comprehensive evaluation across different model families limits the practical applicability of these methods.
@@ 4 Challenges and Future Directions
Despite the efforts of metaTextGrad Xu et al. (2024a), which applies meta-learning to automatically optimize templates for evaluator, human intervention remains to craft the meta-learner’s prompts. To move toward truly automated system optimization, akin to near hyperparameter-free neural network training, we urge future research to reduce reliance on both textual and numerical hyperparameters. For any hyperparameters that remain, we advocate thorough sensitivity analyses to help users understand each method’s behavior and robustness.
@@
Since compound AI systems are expected to address complex problems, it is important to investigate their effectiveness on more challenging tasks. However, most of the papers included in this survey primarily evaluate their proposed methods on common, widely used datasets, such as those for math reasoning (e.g., GSM8K Cobbe et al. (2021)), commonsense reasoning (e.g., MMLU Hendrycks et al. (2021)), and code generation (e.g., HumanEval Chen et al. (2021) and MBPP Austin et al. (2021)). While these evaluations reflect general effectiveness, we argue that it is also important to include benchmarks involving more complex tasks. For instance, AgentBench Liu et al. (2023) and AgentGym Xi et al. (2024) comprise multiple tasks requiring different LLMs within the system to cooperate and discuss, and GAIA Mialon et al. (2023) examines system effectiveness in real-world scenarios requiring various tool usage. Despite the existence of these benchmarks, we propose that developing more collaborative or multimodal tasks specifically designed to benchmark AI system optimization methods is a promising direction for future research.
@@
While NL Feedback methods have shown promising empirical results, they lack theoretical guarantees. For example, the convergence of textual gradient descent remains unproven, whereas classical gradient descent are supported by formal convergence proofs Cheridito et al. (2022); Hutzenthaler et al. (2021). Such proofs provide a solid foundation for the continuous advancement of single-model optimization. We therefore advocate future work to deliver rigorous convergence and optimality analyses for learning via NL Feedback, which will offer deeper insights and strengthen the field’s theoretical underpinnings.
@@ 5 Conclusions
We survey recent advances in optimizing compound AI systems composed of interacting components like agents and tools. To unify diverse research efforts, we propose a graph-based formalism with conditional edges, enabling structured analysis of system interactions. With this framework, we examine existing methods across four principled dimensions and organize them into a 2$\times$2 taxonomy. Our survey highlights key trends and trade-offs, including computational overhead, the NL interface, and challenges in scaling and generalization. We also identify open problems and outline future directions to guide continued progress in the field.
@@
Second, we focus exclusively on methods that explicitly optimize systems of multiple nodes, thereby excluding traditional prompt optimization techniques for standalone LLMs and contributions not framed as system optimization. Despite our efforts to draw a clear boundary, some relevant papers may have been inadvertently omitted; moreover, as this field is actively evolving during the preparation of this manuscript, studies published in the last two months may be only partially covered. To address this, we maintain an open-source repository, released upon publication, where readers can submit works we may have missed.
@@
Appendix A More Details of Formal Definitions

### v2

@@ 1 Introduction
The community has witnessed a new generation of AI systems centered on large language models (LLMs), incorporating several sophisticated components such as simulators, code interpreters, web search tools, and retrieval-augmented generation (RAG) modules. These systems have shown remarkable capabilities across domains and typically outperform standalone LLMs. For instance, LLMs communicating with symbolic solvers can tackle Olympiad-level math problems Trinh et al. (2024), integrate with search engines and code interpreters to match human programmers’ performance Li et al. (2022); Yang et al. (2024b), and when coupled with knowledge graphs, drive biological materials discovery Ghafarollahi and Buehler (2024).
Figure 1: The proposed $\times2$ taxonomy spans Structural Flexibility (y-axis) and Learning Signals (x-axis). Representative methods for each quadrant along with their key designs and potential drawbacks.
@@
However, even with mature toolkits that streamline the design process of compound AI systems, such as, LangChain Chase (2022) and LlamaIndex Liu (2022), substantial human intervention remains essential to tailor these systems toward targeted downstream applications Xia et al. (2024); Zhang et al. (2024c), often involving trial-and-error tuning prompt templates and system pipelines based on heuristics. This limitation has motivated recent efforts to develop principled, automated methods for end-to-end AI system optimization. Yet, the operational schemes of these approaches diverge significantly depending on whether they permit modifications to the system topology and how they transmit learning signals. Moreover, the field still lacks standardized terminology or a cohesive conceptual framework, making some articles difficult for newcomers to navigate Cheng et al. (2024); Khattab et al. (2023). Despite existing surveys Lin et al. (2024); Liu et al. (2025) have provided helpful frameworks, they focus on optimization driven by natural language, overlook critical schemes that permit updates to system topology and do not cover the most recent advances.
To address these gaps, this study introduces four key dimensions for examining existing methods and, drawing on two of these dimensions, constructs a $\times2$ taxonomy (Fig. 1) that holistically covers 26 representative works. Setting a reasonable scope for our survey, we excluded related but not directly relevant work. These include prompt optimization techniques for a single LLM Zhou et al. (2022b); Pryzant et al. (2023); Yang et al. (2023); Das et al. (2024); Guo et al. (2025), systems built for task-specific applications without optimization Zhuge et al. (2023); Hong et al. (2023), and papers that do not explicitly frame their problem as system optimization Zhou et al. (2022a); Sordoni et al. (2023). The proposed taxonomy is inherently adaptable to future research, and we hope it will enable developers of compound AI systems to gain a high-level overview before delving into technical details, while allowing researchers proposing novel algorithms to better situate and compare their contributions.
The remainder of this paper is organized as follows: Sec. 2 provides the necessary background and formalization for compound AI systems and their optimization. Sec. 3.1 discusses the four key dimensions, followed by Sec. 3.2, which presents a detailed review of surveyed papers based on the proposed $\times2$ taxonomy. Sec. 4 identifies the main challenges facing current methods and their corresponding future directions, and Sec. 5 concludes with a summary of our contributions.
@@ 2.2 Formal Definitions
Specifically, we use $\Phi=(G,\mathcal{F})$ to denote a compound AI system that inputs a user query $q\in\mathcal{Q}$ and outputs a response (answer) $a\in\mathcal{A}$. In particular, $G=(V,E)$ is a directed graph and $\mathcal{F}=\{f_{i}\}_{i=1}^{|V|}$ is a set of operations (e.g., LLM forward pass, RAG step, external tool calling). Each $f_{i}$ is attached to node $v_{i}$, producing
@@ 3.2 Representative Methods
Table 1: Compound AI System Optimization methods, sorted by their first publication date on arXiv. All methods and their properties along the four principled dimensions are listed. Superscripts a, $^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{b_{1}}}}$, $^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{b_{2}}}}$, and $^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\bm{b_{3}}}}$ denote the type of numerical signal each method employs (Fig. 3). An asterisk (*) indicates methods that require model fine-tuning. For graph-based system representations, a dagger ($\dagger$}}) marks methods restricted to acyclic structures (i.e., DAGs).
@@
Serving as pioneers in the field, these methods demonstrate that learning from text is possible not only for single AI models but also for systems composed of discrete modules. Nevertheless, their successes are supported mainly by empirical results without theoretical grounding. In addition, the common reliance on proprietary LLMs to generate NL feedback incurs high API costs.
@@
Rather than relying on textual signals, methods in this category use numerical signals to update node parameters. DSPy Khattab et al. (2023) provides a Python library [^5] http://dspy.ai featuring declarative programming modules for designing and optimizing $\Phi$. As a method that learns directly from raw system metrics, it introduces a suite of rejection‐sampling–based routines (Bootstrap-*) for generating high‐quality in‐context demonstrations informed by corresponding system performance. Users may optionally fine‐tune LLM weights ($\theta_{i,N}$) on the collected demonstrations via BootstrapFinetune procedure. MIPRO Opsahl-Ong et al. (2024) advances by jointly optimizing demonstrations and instructions. Specifically, it employs a Bayesian surrogate model to maintain and update posterior distributions over instruction–demonstration configurations, favoring those that yield high performance. BetterTogether Soylu et al. (2024) extends by alternating between LLM prompt and weight optimization, enabling LLMs to iteratively teach themselves and outperform single-strategy baselines.
@@
These methods introduce novel ways to leverage system performance signals, effectively mitigating the imprecision of natural language. However, excessive human‐designed rules may limit generalizability. Additionally, fine‐tuning each LLM in $\Phi$ incurs substantial GPU resource demands.
@@
Methods discussed previously tune only node parameters within a predefined system topology. To overcome this constraints, methods in this category leverage NL Feedback to jointly optimize both node parameters and overall system structure. Concurrent with TextGrad Yuksekgonul et al. (2025), Agent Symbolic Learning Zhou et al. (2024) designs optimizers with three components: PromptOptimizerToolOptimizer, and PipelineOptimizer, making it goes beyond node tuning because the latter two components support tool creation as well as node addition, deletion, and movement. Through a series of analytical experiments, MASS Zhou et al. (2025) demonstrates that optimizing LLM prompts $\{\theta_{i,T}\}$ can deliver performance gains in compound AI systems more easily than exploring the topology $(V,E)$. Building on this insight, MASS devises a three‐stage framework in which prompt optimization precedes topology search.
@@
Methods in this category demonstrate how novel algorithms can leverage the power of proprietary LLMs for direct workflow design. Nonetheless, effective engagement with these meta LLMs demands high token consumption Hu et al. (2024) or multiple LLM inferences Su et al. (2025). Also, evaluations on less powerful open-source models remain scarce.
@@
Completing our review, this passage describes methods that employ Numerical Signals to refine $\Phi$ without posing fixed topology constraints. DyLAN Liu et al. (2024) and GPTSwarm Zhuge et al. (2024) are methods in this category that do not involve model fine-tuning. DyLAN models multiturn debate as a temporal feed-forward network in which agents prompted with distinct roles are unrolled across layers. The system is optimized through pruning unhelpful agents and adaptively connect the surviving agents between consecutive layers. Rule-based algorithms are employed to score each agent’s importance, positioning DyLAN within the category that leverages raw system metrics as learning signals. GPTSwarm models $\Phi$ as a hierarchical framework composed of nodes ($v_{i}\in V$), agents (graphs that link nodes), and swarms (composite graphs that interconnect multiple agents), then employs an optimization procedure to refine the edge connections among agents. A parameterized probabilistic distribution $D_{\theta}$ is introduced to govern the connectivity within the swarm graph, which is then optimized using a gradient‐ascent variant of the REINFORCE Williams (1992) algorithm, placing GPTSwarm within the class of RL methods.
Before discussing the remaining methods, we note that they operate in a query‐adaptive manner, i.e., instantiating a new $\Phi$ for each $q_{i}$. This operating scheme contrasts with methods discussed so far in which a single task‐specific $\Phi$, once optimized, is used for all $q_{i}$. Similar to ADAS Hu et al. (2024) and AFlow Zhang et al. (2024a), these methods leverage a meta LLM to generate optimal system designs. However, rather than relying solely on inference, they fine-tune the meta LLM using constructed training datasets or reward functions to enable effective pipeline generation, given the limited prior knowledge about compound AI system design within LLMs
Starting from the SFT category, MAS-GPT Ye et al. (2025) first constructs a query pool from open‐source queries across various domains, and a MAS pool by manually implementing over 40 common system designs in Python code. Through evaluation, selection, and refinement of query–MAS pairs, MAS-GPT curates a training dataset that ensures consistency, minimizing ambiguity when fine‐tuning the meta LLM.
@@
Recent trends in this category frame compound system optimization as a meta LLM fine‐tuning problem, thereby sidestepping the non‐differentiability of modules within $\Phi$. However, substantial effort is still required to curate high‐quality training data, and the lack of comprehensive evaluation across different model families limits the practical applicability of these methods.
@@ 4 Challenges and Future Directions
Despite the efforts of metaTextGrad Xu et al. (2024a), which applies meta‐learning to automatically optimize templates for evaluator, human intervention remains to craft the meta‐learner’s prompts. To move toward truly automated system optimization, akin to near hyperparameter‐free neural network training, we urge future research to reduce reliance on both textual and numerical hyperparameters. For any hyperparameters that remain, we advocate thorough sensitivity analyses to help users understand each method’s behavior and robustness.
Figure 4: The frequency statistics of benchmarks tested in the surveyed 26 papers.
@@
Since compound AI systems are expected to address complex problems, it is important to investigate their effectiveness on more challenging tasks. Yet, papers in the field primarily evaluate their proposed methods on datasets widely used for single LLMs (Fig. 4), such as those for math reasoning (e.g., GSM8K Cobbe et al. (2021)), commonsense reasoning (e.g., MMLU Hendrycks et al. (2021)), and code generation (e.g., HumanEval Chen et al. (2021) and MBPP Austin et al. (2021)). While these evaluations reflect general effectiveness, we argue that it is also important to include benchmarks involving more complex tasks. For instance, AgentBench Liu et al. (2023) and AgentGym Xi et al. (2024) comprise multiple tasks requiring different LLMs within the system to cooperate and discuss, and GAIA Mialon et al. (2023) examines system effectiveness in real-world scenarios requiring various tool usage. Furthermore, given the broad utility of compound AI systems (e.g., healthcare systems with doctors embedded as nodes in $\Phi$ Chen et al. (2024)), it is necessary to evaluate algorithmic performance when humans function as nodes within the system and to devise principled methods for modeling their behavior.
@@
While NL Feedback methods have shown promising empirical results, they lack theoretical guarantees. For example, the convergence of textual gradient descent remains unproven, whereas classical gradient descent are supported by formal convergence proofs Cheridito et al. (2022); Hutzenthaler et al. (2021). Such proofs provide a solid foundation for the continuous advancement of single‐model optimization. We therefore advocate future work to deliver rigorous convergence and optimality analyses for learning via NL Feedback, which will offer deeper insights and strengthen the field’s theoretical underpinnings.
Potential Safety Risks
While safety issues and corresponding defenses for single models have been extensively studied, such as jailbreak attacks on LLMs and their mitigation Yi et al. (2024) and human-engineered AI pipelines to reduce harmful prompts and outputs Han et al. (2024), the attack surface expands considerably in compound AI systems Banerjee et al. (2024). For instance, Debenedetti et al. (2024) demonstrate that privacy-preserving models can still leak sensitive data when integrated as a component of a larger system. Furthermore, because compound AI systems are often embodied and executed as code within enterprise environments, latent failure modes may remain undetected and undermine system reliability even in the absence of explicit adversarial attacks. Despite these risks, research on compound AI system optimization that extends beyond downstream performance has so far addressed mainly execution efficiency Zhang et al. (2025), with little attention to system-level alignment or safety Zheng et al. (2025). Given the mature alignment and safeguarding optimization techniques available for single models Achiam et al. (2017); Dai et al. (2023), we urge future work to adapt and extend these strategies to compound systems in order to balance capability enhancements with safety guarantees Yang et al. (2024a).
@@ 5 Conclusions
We survey recent advances in optimizing compound AI systems composed of interacting components like agents and tools. To unify diverse research efforts, we propose a graph-based formalism with conditional edges, enabling structured analysis of system interactions. With this framework, we examine existing methods across four principled dimensions and organize them into a $\times2$ taxonomy. Our survey highlights key trends and trade-offs, including computational overhead, the NL interface, and challenges in scaling and generalization. We also identify open problems and outline future directions to guide continued progress in the field.
@@
Second, we focus exclusively on methods that explicitly optimize systems of multiple nodes, thereby excluding traditional prompt optimization techniques for standalone LLMs and contributions not framed as system optimization. Despite our efforts to draw a clear boundary, some relevant papers may have been inadvertently omitted; moreover, as this field is actively evolving during the preparation of this manuscript, studies published in the last two months may be only partially covered. To address this, we maintain an open‐source repository where readers can submit works we may have missed.
@@
Acknowledgements
We thank the reviewers for their insightful comments. This work was financially supported by the National Science and Technology Council (NSTC) in Taiwan, under Grants 111-2222-E-002-013-MY3 and 112-2223-E002-012-MY5, and the Center of Data Intelligence: Technologies, Applications, and Systems, National Taiwan University under the Grant 114L900901, from the Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE) of Taiwan.
Appendix A More Details on Formal Definitions